a hollow metal sphere of mass m has an instantaneous upward acceleration a when released from rest fully submerged in a liquid of density d1.if the density of the metal is d2 find the volume of cavity in the sphere 

Grade:12th pass

1 Answers

Wonder
6 years ago
Draw FBD; F (b)-mg=ma total volume×d1×g-volume(metal sphere)d2g=V(metal sph)d2×a.write total volume in form of volume of metal+cavity.{V (m)+V (c)}d1g=V (m)d2 (g+a). d2×V(m)+V (c)d2÷d2=V (m)d2 (g+a)/d1g. Here V (m)d2 is mass of sphere -m. So m÷d2+V(c)=m (g+a)/d1g;; V(c)=m (g+a)/d1g-m/d2.